A body of mass 40 kg stands on a weighing machine in a accelerated lift. The reading on the scale of the weighing machine is 300N. Find the magnitude and direction of acceleration. (g=9.8 m/s^2)

Grade:9

1 Answers

Arun
6 years ago
Dear student
 
M=40kg
F=300N
suppose that g=9.8m/s²

Now, when the lift goes down, 2 types of accelerations are applied, 
1. gravitational acceleration
2.acceleration of life

By thinking about the FBD of the lift,
if the lift moves down, the sum of acceleration = g-a
and if it moves up, it is g+a

[why? Becuase when the life moves, an illusory acceleration is applied on it, which is 'a' here]

That means, 
1. if the lift is moving upwards,
F=m(g+a)
300=40(10+a)
a= -2/5 m/s²

2. if lift is going down,
F=m(g-a)
300=40(10-a)
a= +2/5 m/s²

Now the directions:
 1.if the lift moves downwards, the magnitude of acceleration is +2/5 m/s²
2.if the lift moves downwards, the magnitude of acceleration is -2/5 m/s²
 
Regards
Arun (askIITians forum expert)